# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

### v1

@@ C.1 GNN training
For all datasets, we implement a GCN with 3 convolutional layers, with the input and output channel dimensions chosen according to each dataset’s feature dimension and number of classes, respectively. For the hidden dimension, we choose either $64$ or $512$, depending on the nodes’ feature dimension. The model is trained with the Adam optimizer Kingma and Ba (2014) and a learning rate of 0.001 for 400 epochs.
C.2 GNN accuracy
We give the GNNs accuracy for the different folds in Table 6.
Table 6: Accuracy of GCN on the node classification task for the Cora, Wiki CS, Liar and Amazon products datasets.
@@ C.2 GNN accuracy
Appendix D Prompt and generated responses examples
In this section we present a detailed prompt construction in D.1 and a complete example of such built prompt with its input in D.2.
D.1 Example of Full Prompt Construction for Logic
We have provided an example of the full prompt used in Logic below. The boldface texts denote where the embeddings are injected into the prompt. An example response is also illustrated in Figure 3 in the paper.
Prompt:
You are an expert research assistant. Your goal is to assess whether the classification of a target product is supported by its co-purchase neighborhood by analyzing Amazon product reviews and its co-purchase neighborhood to understand why it has been classified under a specific category. The categories to distinguish from are Home & Kitchen, Health & Personal Care, Beauty, etc.
Target Product ID: 0
Predicted Category: Clothing, Shoes & Jewelry
Target Product Embedding Representation:
[…]
Neighboring Products in the Co-purchase Network:
@@ D.1 Example of Full Prompt Construction for Logic
  • Product 129:
[…]
  • Product 130:
[…]
Instructions:
You are given a target product and a list of neighboring products, each described by keywords.
For each neighboring product:
  1. 1.
    Write one sentence summarizing the main topics or ideas captured in its keywords.
  2. 2.
    Clearly state whether this product supports the classification of the target product into category ‘Clothing, Shoes & Jewelry’.
Use the following format for each neighbor:
Product {ID}:
Summary: One sentence summary of the product’s keywords.
@@
Support: YES or NO — Does this product support the classification into ‘Clothing, Shoes & Jewelry’?
Base your reasoning only on the keywords and proximity to the target product.
Start your analysis below:
D.2 Sample of the prompt and output for the Amazon-Product dataset
We provide a sample of the prompt and a sum up of the output we obtain (without the supporting sentence for each product) in Figure 3.
Figure 3: An illustration of the prompting template and LLM explanation of Logic for the Amazon dataset. The left block shows the input prompt to the LLM for a sample target node (target node id 3). The right block denotes the LLM response explaining the prediction of the GNN for the target node.
Appendix E Additional Experiments
We extend Table 1 of the paper with additional experiments for Amazon with TAGExplainer Pan et al. (2024), PGExplainer Luo et al. (2020) and with the GIN Xu et al. and GAT Veličković et al. (2017) architectures. The results are reported in Table 7.
Table 7: Performance comparison on the Products dataset using different GNN architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
GCN GAT GIN
Fidelity Size Fidelity Size Fidelity Size
Node 73.8% 1.00 73.2% 1.00 39.6% 1.00
Random 84.4% 7.24 84.2% 7.28 90.2% 7.25
GNNExplainer 95.5% 5.63 94.5% 5.615 88.0% 5.65
PGExplainer 87.2% 5.46 88.3% 5.76 87.2% 5.46
TAGExplainer 73.8% 1.37 71.8% 1.43 42.3% 1.27
LLM 86.5% 5.07 91.0% 5.00 86.5% 5.18
Our method 87.0% 1.3 81.5% 1.45 60.0% 1.52
Our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance for the GCN with 87.0% fidelity at just 1.3 in size. However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (60.0%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation. In contrast, GNNExplainer maintains the highest fidelity across all architectures but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes. The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise. These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structural-based models.
Fidelity Size Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 86.5% 5.07 88.0% 5.14 84.0% 2.33 88.0% 5.17 84.0% 2.34
Our method 87.0% 1.30 83.0% 1.63 80.0% 1.00 80.0% 1.00 80.0% 1.00
The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger. In contrast, our method consistently generates more compact explanations—typically around size 1, with a moderate decrease in fidelity of roughly 4–7 percentage points depending on the model. Notably, Llama 3.1 and Mistral backbones show strong fidelity for both methods, whereas smaller models like GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall. These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
Appendix F Definition of the qualitative metrics
@@
Appendix G Experimental Setup for the human evaluation of M1 and M2 Scores
We clarify the experimental setup used to obtain the M1 and M2 scores reported in Table 2.
@@ Appendix G Experimental Setup for the human evaluation of M1 and M2 Scores
We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained. For each article, we generated explanations using both LOGIC and GNNExplainer. Each expert was provided with:
@@
  • Before the evaluation, we provided experts with definitions of the evaluation metrics (as described in Appendix F) and asked them to consider how to apply each metric to content from the Cora dataset.
@@
The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Logic. These results are reported in Table 5.3.

### v3

@@ D.1 Experimental Setup
We use a GCN Kipf and Welling (2017b) as the base GNN model for all datasets. The GCN consists of 3 convolutional layers, with the final layer used for classification. The model is trained for 100 to 400 epochs (tune based on validation set for each dataset) using the Adam optimizer with a batch size of 64. For the hidden dimension, we choose either $64$ or $512$, depending on the nodes’ feature dimension.
We provide the GNN’s accuracy where GCN is used as the model architecture for the different folds in Table 11.
@@
Table 11: Accuracy of GCN for the node classification task on four datasets.
For main experiments, we use Meta’s Llama-3.1-8B-Instruct model from the HuggingFace library as a frozen LLM.
Projector.
The projector is implemented as a two-layer MLP with ReLU activation. The hidden dimension is set to 4x the GNN embedding dimension. The output layer maps the GNN embedding to a vector of size $k\cdot h$, where $k$ is the number of soft prompt tokens and $h$ is the LLM embedding dimension. The mean-pooled representation is used for alignment objectives. We fix $k=5$ across all datasets.
The projector is trained separately for each dataset using full-batch gradient descent. We use the Adam optimizer, with the learning rate selected via random search over $\{0.001,0.0005\}$ based on validation performance. The number of training epochs is selected from $\{100,200,400\}$ using validation. The loss weights are selected via random search over $\{0.1,0.5,0.9\}$ on the validation set. The contrastive temperature is fixed to $\tau=0.2$.
Explanation construction.
For each node, we construct explanation subgraphs by including all supporting nodes and excluding all opposing nodes.
LLM inference.
We use the default HuggingFace generation settings for the LLM, without modifying temperature or sampling parameters, and set the maximum number of generated tokens to 1024. Prompts are formatted using the chat template when supported by the model, and a fixed prompt template is used across all datasets.
Baselines and evaluation protocol.
We compare Gspell against standard GNN explainers including GNNExplainer, PGExplainer, and TAGExplainer. We use official implementations and PyTorch Geometric implementations where applicable, without additional hyperparameter tuning.
All methods are evaluated under the same GNN backbone for fairness. Experiments are repeated over 5 random seeds, and we report the mean and standard deviation of all metrics.
We discuss additional details related to the training of the base GNN.
D.2 Additional example on Amazon-Product
A key advantage of Gspell is its ability to increase human understandability and trustworthiness by providing explanation of explanations. We give an additional example on Amazon-Product in Figure 3. The “Reasoning” section of the LLM response is where the LLM provides a descriptive reasoning on how the supporting neighbors can explain the prediction of the target node. On the other hand, the previous GNN explainers simply generate a subgraph that is more abstract and harder to understand without some level of background knowledge.
LLM Prompt You are analyzing Amazon product reviews and their co-purchase neighborhoods to understand why it has been classified under a specific category. Your task is to determine which of the neighboring products does (not) support the classification of the target product. Target Product ID: 160 Predicted Category: Health & Personal Care
Product Embedding Representation:
<<BEGIN TARGET KEYWORDS>> PROJECTED EMBEDDING <<END TARGET KEYWORDS>>
Neighboring Products:
@@ D.2 Additional example on Amazon-Product
Product 200: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 201: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 202: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 203: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 204: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 205: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 206: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 207: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 208: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 209: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 210: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
LLM Response P200 Summary: Home care and hygiene products. NO P201 Summary: Personal grooming and beauty tools. NO P202 Summary: Health and wellness supplements. YES P203 Summary: Fitness and exercise equipment. NO P204 Summary: Skincare and beauty products. YES P205 Summary: Health and wellness journals. YES P206 Summary: Personal care and hygiene for specific needs. YES P207 Summary: Health and wellness supplements. YES P208 Summary: Fitness and exercise accessories. NO P209 Summary: Health and wellness products for specific needs. YES P210 Summary: Health and wellness journals. YES Reasoning: The common pattern among the supporting Products is that they are all related to health and wellness, skincare, and personal care, which are all categories within the Health & Personal Care category.
Instructions For each neighboring product: 1. Summarize the similarity between supporting products in one sentence. 2. Clearly state whether this product supports the classification of the Target Product into category Health & Personal Care. Format:
Product <ID>:
Summary: <One sentence>.
@@
Support: YES or NO — Does this product support classification into Health & Personal Care?
Base reasoning only on the keywords and proximity to the target product.
Figure 3: Left: prompt with category and embeddings highlighted. Right: model response with summaries, YES/NO verdicts, and reasoning. Below: instructions shown separately.
Appendix E Definition of the qualitative metrics
@@
Appendix F Experimental Setup for the human evaluation of M1 and M2 Scores
The human evaluation was conducted by the authors, who served as domain experts. For the qualitative experiments, one expert reviewed the Cora dataset, and four experts reviewed the Amazon-Product data, evaluating the explanations generated by GNNexplainer and Gspell over 10 node classification explanation tasks. We clarify the experimental setup used to obtain the M1 and M2 scores reported in Table 2.
@@ Appendix F Experimental Setup for the human evaluation of M1 and M2 Scores
We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained. For each article, we generated explanations using both Gspell and GNNExplainer. To mitigate potential bias, explanations from these methods were randomized and presented in a blind manner, without revealing their source. Each expert was provided with:
@@
  • Before the evaluation, we provided 3 experts with the evaluation metrics’ definitions and asked them to consider how to apply each metric to content from the Cora dataset.
@@
The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Gspell. These results are reported in Table 2.